# Actions Speak Louder Than Words: Rate-Reward Trade-off in Markov Decision Processes

### v2

@@ C.5 Vector Representation of States and Actions
To evaluate the effectiveness of Act2Comm with vector representations, we applied this variant to the “Catch the Ball” scenario with $R=1/3$ (cf. Fig. 4(a)). In this environment, the state is represented as a vector $[s_{a},s_{b}]$, where $s_{a}\in\{0,1,2\}$ indicates the position of the board, and $s_{b}\in\{0,1,2,3,4,5,6,7,8\}$ indicates the position of the ball. The performance of Act2Comm with vector representations in this experiment is very similar to that achieved with scalar state representations, as shown in Fig. 4. Notably, we observe that the vector representation offers advantages such as more stable and faster convergence during training. This is illustrated in Fig. 8(a), where the loss value decreases and converges more quickly during training compared to the scalar state representation counterpart. An intuitive explanation for this result is that the vector state representation captures more detailed features of the environment, which may potentially facilitate learning.
@@ D.2 Lucky Wheel
As illustrated in Fig. 9(a), the wheel is evenly divided into three regions. At each time step, the player can select either action 00 or action $1$, where action 00 corresponds to a clockwise rotation of the wheel, and action $1$ corresponds to an anti-clockwise rotation. If action 00 is selected, there is a probability $p=0.2$ that the pointer remains in its current region, and a probability $0.81-p=$ 0.8 that it moves to the next region in the clockwise direction. Similarly, if action $1$ is selected, there is a probability $p=0.2$ that the pointer remains in its current region, and a probability $0.81-p=$ 0.8 that it moves to the next region in the anti-clockwise direction. The rewards for the three regions are $5$, $-5$, and 00, respectively. The player receives a reward at each time step based on the region in which the pointer is located.
@@ D.3 Catch the Ball
The “Catch the Ball” game is set in a $3\times 3$ grid, as illustrated in Fig. 9(b). A ball randomly appears at the top of the grid and descends one grid space at each time step. Meanwhile, a board at the bottom moves horizontally to catch the falling ball. Each time the board successfully catches a ball, the player receives a reward $r$. If the ball falls to the bottom of the grid without being caught, it disappears, resulting in a penalty of $-r$ for the player. At all other times, there are no rewards or penalties for the player. After a ball disappears or is caught, a new ball appears randomly (with equal probability) at one of the three positions at the top of the grid.
The player has three available actions to move the board: move left, move right, or remain stationary. If the player chooses to move left or right, there is a probability $p\in[0,1)$ that the movement fails (in which case the board remains in its current position), and a probability $1-p$ that the board moves for one grid space successfully as intended. Additionally, any action that attempts to move the board outside the grid will always fail. In this experiment, we set $r=5$ and $p=0.8$.
@@
The ‘Erratic robot’ game takes place on a $4\times 4$ grid map with $16$ states and $5$ actions, as shown in Fig. 9(c). The robot is designed to minimize the number of steps required to collect goods while avoiding obstacle points. Specifically, its primary objective is to continuously take goods from designated destination points, earning a reward of $+5$ for each successful collection. The grid also includes four obstacle points, each incurring a penalty of $-2$ when encountered. The robot has five available actions: move left, move right, move up, move down, or remain stationary. Due to the instability, any movement of the robot has a probability $p\in[0,1)$ of resulting in an additional unintended step. Actions that would move the robot outside the grid boundaries always fail. In this experiment, the probability of additional operation is set to $p=0.2$. The results are presented in the Fig. 10.
@@ Appendix E Additional ablation experiments
This section presents additional ablation experiments for our proposed Act2Comm framework, using the lucky wheel environment as the default experimental setting with an initial state 00.

### v3

@@ C.5 Vector Representation of States and Actions
To evaluate the effectiveness of Act2Comm with vector representations, we applied this variant to the “Catch the Ball” scenario with $R=1/3$ (cf. Fig. 4(a)). In this environment, the state is represented as a vector $s_{a},s_{b}]$, where $s_{a}\in\{0,1,2\}$ indicates the position of the board, and $s_{b}\in\{0,1,2,3,4,5,6,7,8\}$ indicates the position of the ball. The performance of Act2Comm with vector representations in this experiment is very similar to that achieved with scalar state representations, as shown in Fig. 4. Notably, we observe that the vector representation offers advantages such as more stable and faster convergence during training. This is illustrated in Fig. 8(a), where the loss value decreases and converges more quickly during training compared to the scalar state representation counterpart. An intuitive explanation for this result is that the vector state representation captures more detailed features of the environment, which may potentially facilitate learning.
@@ D.2 Lucky Wheel
As illustrated in Fig. 9(a), the wheel is evenly divided into three regions. At each time step, the player can select either action 0 or action $1$, where action 0 corresponds to a clockwise rotation of the wheel, and action $1$ corresponds to an anti-clockwise rotation. If action 0 is selected, there is a probability $p=0.2$ that the pointer remains in its current region, and a probability 1p=0.81-p=0.8 that it moves to the next region in the clockwise direction. Similarly, if action $1$ is selected, there is a probability $p=0.2$ that the pointer remains in its current region, and a probability 1p=0.81-p=0.8 that it moves to the next region in the anti-clockwise direction. The rewards for the three regions are $5$, 5-5, and 0, respectively. The player receives a reward at each time step based on the region in which the pointer is located.
@@ D.3 Catch the Ball
The “Catch the Ball” game is set in a $3\times 3$ grid, as illustrated in Fig. 9(b). A ball randomly appears at the top of the grid and descends one grid space at each time step. Meanwhile, a board at the bottom moves horizontally to catch the falling ball. Each time the board successfully catches a ball, the player receives a reward $r$. If the ball falls to the bottom of the grid without being caught, it disappears, resulting in a penalty of r-r for the player. At all other times, there are no rewards or penalties for the player. After a ball disappears or is caught, a new ball appears randomly (with equal probability) at one of the three positions at the top of the grid.
The player has three available actions to move the board: move left, move right, or remain stationary. If the player chooses to move left or right, there is a probability $p\in[0,1)$ that the movement fails (in which case the board remains in its current position), and a probability 1p1-p that the board moves for one grid space successfully as intended. Additionally, any action that attempts to move the board outside the grid will always fail. In this experiment, we set $r=5$ and $p=0.8$.
@@
The ‘Erratic robot’ game takes place on a $4\times 4$ grid map with $16$ states and $5$ actions, as shown in Fig. 9(c). The robot is designed to minimize the number of steps required to collect goods while avoiding obstacle points. Specifically, its primary objective is to continuously take goods from designated destination points, earning a reward of +5+5 for each successful collection. The grid also includes four obstacle points, each incurring a penalty of 2-2 when encountered. The robot has five available actions: move left, move right, move up, move down, or remain stationary. Due to the instability, any movement of the robot has a probability $p\in[0,1)$ of resulting in an additional unintended step. Actions that would move the robot outside the grid boundaries always fail. In this experiment, the probability of additional operation is set to $p=0.2$. The results are presented in the Fig. 10.
@@ Appendix E Additional ablation experiments
This section presents additional ablation experiments for our proposed Act2Comm framework, using the lucky wheel environment as the default experimental setting with an initial state 0.